OpenStudy (anonymous):

how is x/ (x^2-1) simplified to 1/(x-1) ?

Nnesha (nnesha):

nope that's not right

Nnesha (nnesha):

again \[\huge\rm a^2 - b^2 = (a+b)(a-b)\] difference of square method x^2-1 can't be x(x-1)

OpenStudy (anonymous):

what about x^3-9/ (x+3) ?

Nnesha (nnesha):

x^3 or x^2 ?

OpenStudy (anonymous):

cubed

Nnesha (nnesha):

okay for difference of square method there should be negative sing and both term suppose to be perfect square root

Nnesha (nnesha):

sign** :D

Nnesha (nnesha):

for that one you have to divide x^3 -9 by x+3

OpenStudy (anonymous):

and how do I do that ?

Nnesha (nnesha):

two ways 1) long division 2)synthetic division which one is easy for you ?

OpenStudy (anonymous):

I have no idea

OpenStudy (anonymous):

what's synthetic division ?

Nnesha (nnesha):

hmmm so haven't ' learned this yet? right :( what's your question divide or simplify for 2nd one

OpenStudy (anonymous):

I want to simplify

Nnesha (nnesha):

\[\huge\rm \frac{ x^3 -9 }{ x+3 }\]its already simplified

Nnesha (nnesha):

hm wait is this ur question \[\huge\rm \frac{ x^3 }{ x+3 } -9\] ??

OpenStudy (anonymous):

no it's the first one

Nnesha (nnesha):

difference of cubes ?know anything about this method ?

OpenStudy (anonymous):

a3 + b3 = (a + b)(a2 – ab + b2)

OpenStudy (anonymous):

but cubert 9 is a decimal

Nnesha (nnesha):

yeah

Nnesha (nnesha):

bec 9 is not a perfect cube root

OpenStudy (anonymous):

yeah thanks anyway

Nnesha (nnesha):

np :)